OpenStudy (anonymous):

Evaluate the limits: 1. as x approaches 1, limit of (x^3 - 1)/(x^2- 1) 2. as x approaches 2, limit of (x^3 - 8)/(2x - 4)

OpenStudy (anonymous):

@hartnn

OpenStudy (sirm3d):

factor \(x-1\) in problem (1) and \(x-2\) in problem (2)

OpenStudy (anonymous):

1. how would u do that (x - 1)(x^2 + 1) = x^3 +x -x^2 -1 it does not work

OpenStudy (sirm3d):

\(x^3-1=(x-1)(x^2+x+1)\)

OpenStudy (sirm3d):

make use of the factorization of the difference of two cubes \[a^3-b^3=(a-b)(a^2+ab+b^2)\]

hartnn (hartnn):

ask, if there are still any doubts....

OpenStudy (shamim):

i wanna try to solve it if u want